OpenStudy (anonymous):

How do I simplify the expression 9^6/9^-3

OpenStudy (anonymous):

you have:\[\frac{9^6}{9^{-3}} = 9^6 . 9^3 = 9^9\] ^_^

OpenStudy (anonymous):

approx = 387420489 :)

OpenStudy (anonymous):

Why don't you multiply 9*9??

OpenStudy (anonymous):

when you divide with same base...you subtract the exponents...you add when you mutliply

OpenStudy (anonymous):

keep base...that is the rule

OpenStudy (anonymous):

no need as long as you can change the sign of the power :) to make your life easier ^_^

OpenStudy (anonymous):

thank you